OpenStudy (anonymous):

I dont get it 2x-4=10 ..... I think I bring down the 4 with addition Please help me below ! The senior class at a local high school is selling candy for their upcoming dance. For each candy bar that is sold, the class earns a profit of $3.50. Their goal is to earn $850. If they had $430 saved prior to the fundraiser, which equation will help the class determine how much candy they must sell? $3.50x + $430 = $850 $430x + $3.50 = $850<----- my choice ! $3.50x + $850 = $430 $850x + $3.50 = $430

OpenStudy (anonymous):

2x - 4 = 10 (you add 4 to both sides) +4 +4 ------------- 2x + 0 = 14 2x = 14 x = 14/2 x = 7 Next question : x = number of candy bars sold 3.50x + 430 = 850 <--answer is A

OpenStudy (anonymous):

If you have any questions, please ask them.I will explain the best I can.